OpenStudy (anonymous):

What is the Greatest Common Factor of x^4 and x^3?

OpenStudy (unklerhaukus):

what does the first term look like after you factor out one \(x\)

OpenStudy (anonymous):

0?

OpenStudy (unklerhaukus):

\[x^4=(x)(x^3)\]

OpenStudy (anonymous):

x^4 would be the answer

OpenStudy (anonymous):

right?

OpenStudy (unklerhaukus):

nope

OpenStudy (unklerhaukus):

i have factored out one \(x\), \(x\) and \(x^3\) are factors of \(x^4\)

OpenStudy (unklerhaukus):

i could have done \(x^4=(x)(x)(x)(x)=xxxx\)

OpenStudy (anonymous):

wait so what am i supposed to do next ?

OpenStudy (unklerhaukus):

any number can be factorized into itself and one \(x=(1)(x)\)

OpenStudy (anonymous):

so its x

OpenStudy (unklerhaukus):

\(x^3=(1)(x^3)\) so we see that \(x^3\) is a factor of boh terms